OpenStudy (anonymous):

let p represent the perimeter of an isosceles triangle that has two equal sides of length A and a third side of length B. The formula for the perimeter is p=2a + b. Determine the length of the equal side of an isosceles triangle having perimeter of 3/4 yard and a third side measuring 1/3 yard.

OpenStudy (jdoe0001):

\( perimeter\implies p=2a+b\qquad \begin{cases} p=\frac{3}{4}\\ b=\frac{1}{3} \end{cases}\implies \cfrac{3}{4}=2a+\cfrac{1}{3} \\ \quad \\ \textit{solve for "a"}\)

OpenStudy (anonymous):

thank you very much